OpenStudy (anonymous):

can u help on these? How much should you invest now at 10% annual interest raate to have $8000 toward the purchase of a car in 5 years if the interest in compounded quaterly?

OpenStudy (amistre64):

A = P(1+r/n)^(nt) ; solve for P A/(P(1+r/n)^(nt)) = P

OpenStudy (amistre64):

that got typoed lol

OpenStudy (amistre64):

A/(1+r/n)^(nt) = P thats better: or with latex we get: \[\frac{A}{(1+\frac{r}{n})^{nt}}=P\]

OpenStudy (amistre64):

in this case: A = 8000 r = .08 t = 5 n = 4, since its compounded quarterly

OpenStudy (amistre64):

r = 10% ... bummer; change that to r=.10 NOT .08

OpenStudy (amistre64):

google does this for us: 8000/(1+.10/4)^(4*5) = 4,882.167... http://www.google.com/search?sourceid=chrome&ie=UTF-8&q=8000%2F(1%2B.10%2F4)^(4*5)